Wuttke Corp. wants to raise $3.6 million via a rights offering. The company currently has 460,000 shares of common stock outstanding that sell for $25 per share. Its underwriter has set a subscription price of $15 per share and will charge the company a spread of 2 percent.

If you currently own 3,000 shares of stock in the company and decide not to participate in the rights offering, how much money can you get by selling your rights? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1
Net proceeds per share = 15 * ( 1 - 2% ) 14.7
Shares to be offered = 3600000 / 14.7 244898
Number of rights needed = 460000 / 244898 1.88
Ex-right price = ( (1.88*25)+(1*15) ) / (1.88+1) 21.53
Value of right = 25 - 21.53 3.47
Proceeds to be received from selling rights = 3.47 * 3000 10410